Solve the equation by completing the square: (k-1)(k+7)=9
Answer by rothauserc(4718)

(k-1)(k+7)=9
k^2 +6k -7 = 9
add 7 to both sides of =
k^2 +6k = 16
divide coefficient of k by 2, 6/2 = 3 and 3^2 = 9
k^2 +6k +9 = 16 +9
(k+3)^2 = 25
take square root of both sides of =
k+3 = square root of 25
k = 5 - 3 = 2
k = -5 - 3 = -8